The concentration of B⁻ in a saturated solution of AB₃ is 1.10 x 10⁻⁵ mg/L. What is the Ksp of AB₃ if the MM of B is 17.0 g/mol? 


1
Expert's answer
2021-01-12T06:58:51-0500

AB3 = A+ + 3B-

1.10 x 10-5 mg/L = 1.10 x 10-8 g/L

[B] = 1.10 x 10-8 / 17.0 = 6.47 x 10-10 (mol/L)

[A+] = 1/3 [B-] = 1/3 x 6.47 x 10-10 = 2.16 x 10-10 (mol/L)

Ksp = [A+][B-]3 = 2.16 x 10-10  x (6.47 x 10-10)3 = 5.84 x 10-38